OpenStudy (anonymous):

4x + 4 = 9x - 36 Medal to the correct answer. Please help. Explane your answer....

OpenStudy (anonymous):

x= 8

OpenStudy (anonymous):

how did u get 8

OpenStudy (anonymous):

4x + 4 = 9x - 36 4x + 4 + 36 = 9x - 36 + 36 4x + 40 = 9x 4x - 4x + 40 = 9x - 4x 40 = 5x x=8

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

No problem